OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. M(0, 6), N(6, 0)

OpenStudy (anonymous):

can someone please help me i am i need of some answers and this is one of the hardest problems i have ever had. i have a high b almost to an low A in my math class. i have taking plenty of notes but my brain just refuses to take in what i am seeing and apply it to the rest of my knowledge

OpenStudy (anonymous):

first you would have to find the slope in order to do that you would y2-y1/x2-x1 0=x1 6=y1 6=x2 0=y2

OpenStudy (anonymous):

Have you tried the y and X intercept form?

Directrix (directrix):

M(0, 6) N(6, 0) Get the slope of the line determined by the two points:

Directrix (directrix):

The slope of line MN is -1. Next up, use the slope-point format to get the equation of the line: Use either point, I'm choosing N(6, 0). The format we will use looks like this: y - y1 = m*(x - x1) where (x1,y1) is the point and m is the slope of the line. y - 0 = (-1)*(x - 6) y = -1*x + 6

Directrix (directrix):

This --> y = -1*x + 6 is the slope-intercept form of the line. Your question calls for --> standard form the equation How does the standard form of a linear equation look? Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. So, take this --> y = -1*x + 6 and write it in the form Ax + By = C where A is a positive integer, and B, and C are integers. @brittneysmith